OpenStudy (anonymous):

how to solve 0.2 : ( 2/7 - 1/6 ) divide 0.32 - 11/125^2/3 ?

jhonyy9 (jhonyy9):

so next divide not is paranthese ?

jhonyy9 (jhonyy9):

like divide ( 0,32 - ...

jhonyy9 (jhonyy9):

hope that you know when there are addition,subtraction,multiplication or division ,with parantheses and without parantheses so what you need doing in the first step ?

jhonyy9 (jhonyy9):

ok i will learn to you so when there are these ALL how i have wrote there up befor so than you need doing firstly the operations inside parantheses begin with multiplication or division and just after these continue with additions or subtractions so after you have eliminated these paranthese so just after this you can continue with multiplication or division and after this with addition or subtraction hope so much that is understandably for you ok ?

jhonyy9 (jhonyy9):

came on i like help you solve this

jhonyy9 (jhonyy9):

how many will be 2/7 -1/6 = ?

jhonyy9 (jhonyy9):

or do you know how many will be 11/125^2/3 ?

jhonyy9 (jhonyy9):

how do you think 125^2/3 =cuberoot(125^2) =cuberoot(5^6) =5^6/3 =5^2 =25 so how do you think this is right ?

jhonyy9 (jhonyy9):

so you dont like cooperating with me bye